# Tuning LLM-based Code Optimization via Meta-Prompting: An Industrial Perspective

### v1

@@ VI-B Actionable Guidelines for Context Provision
VI-C Actionable Guidelines for Meta-Prompting LLM Selection
@@
VI-D Limitations and Future Work
@@ VII Conclusion
This paper addresses the cross-model prompt engineering bottleneck in industrial LLM-based code optimization. Across 366 hours of benchmarking, we show that MPCO achieves significant performance improvements across five real-world codebases while preserving industrial efficiency constraints. This enables platforms like Artemis to adopt multi-LLM optimization strategies without manual prompt engineering overhead. Through detailed analysis, we provide actionable recommendations for industrial deployment:
Maintain comprehensive context integration.
Automate context collection using existing infrastructure.
Select meta-prompting LLMs based on deployment constraints rather than minor LLM performance differences.
Consider advanced meta-prompting techniques when trading off computational cost for enhanced performance.
Future work of this study includes extending MPCO to handle multi-level and multi-objective optimization scenarios.

### v2

@@ VI-D Actionable Guidelines for Context Provision
VI-E Actionable Guidelines for Meta-Prompting LLM Selection
@@
VI-F Limitations and Future Work
@@
Verification reliability: Our validation of functional correctness relies on existing unit test suites of target projects—if original test coverage is low, LLM-generated optimized code might introduce subtle bugs or edge case failures that go undetected. Future work should investigate integrating automatic test case generation, property-based testing, or symbolic checks to create more robust verification pipelines that are less dependent on existing test coverage.
@@ VII Conclusion
We have presented Mpco, a meta-prompting approach that automatically optimizes and selects effective prompts for code optimization across multiple LLMs. Over 366 hours of benchmarking, Mpco demonstrated significant performance gains on five real-world codebases while meeting industrial efficiency requirements. This makes it practical for automated code optimization platforms such as Artemis to adopt multi-LLM optimization strategies without the burden of manual prompt engineering.
Based on our analysis, we recommend maintaining comprehensive context integration, automating context collection using existing infrastructure, selecting meta-prompting LLMs according to deployment constraints rather than marginal model differences, and applying advanced techniques only when higher performance justifies the additional computational cost. Our future work will extend Mpco to support multi-level and multi-objective optimization scenarios, as well as integrating automatic test generation for more robust verification pipelines.
Acknowledgment
This work was supported in part by an Innovative UK Knowledge Transfer Partnership (KTP) project between the University of Leeds and TurinTech AI, and the UK Engineering and Physical Sciences Research Council (EPSRC) under grant agreement EP/X037304/1.